OpenStudy (anonymous):

Algebraic refresh help needed... The x-intercepts of y= -1/2 cos (x/2), are, pi/2 & 3pi/2, but "HOW" algebraically do I find these points?.... I know you set y=0 and so cos of 0 is 1 which is pi/2, but I don't know how "algebraically" to find 3pi/2?... I know this is a simple explanation but im stuck

OpenStudy (anonymous):

thank you anyways

OpenStudy (anonymous):

So the problem is... \[y =\frac{ -1 }{ 2 } \cos \frac{ X }{ 2 }\] and the x-intercepts are \[\frac{ 3\Pi }{ 2 }\ and \frac{ \Pi }{ 2 }\] You can try working backwards. I'm not sure this is how your teacher wants it but this is how I worked it out. Instead of putting Y as 0, set X equal to its intercept. I wrote this on a sticky note so excuse the lame picture, I didn't want to re-write everything on here. (SORRY FOR HOW LONG THIS TOOK!)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome and good luck! I'm working through trigonometry right now and all these sines and cosines and tangents and their inverses are getting all muddled up in my head! :p

OpenStudy (anonymous):

I have a similar question thats y= -2 cos x/4 , and i know to find the period it's 2pi/b but in this case is it, 2pi/4 or 2pi/1? i thought it was 2pi/1?